4x+9x-6x=42 what does x equal and show work. thank you
Answer by tutorcecilia(2152)

4x+9x-6x=42
13x-6x=42
7x=42
7x/7=42/7
x=6
.
Check by plugging (x=6) back into the original equation.
4x+9x-6x=42
4(6)+9(6)-6(6)=42
42=42 [Checks out]